Grade 12Algebra

If G1 and G2 are the two G.M.s between b,c and A is their A.M., then show that G1^3 + G2^3=2Abc.

Ali Raza
8 Years agoGrade 12

1 Answer

Samyak Jain
8 Years ago
G1 and G2 are the two G.M.s between b,c and A is their A.M.
b   G1   G2   c .  By using basic definition of G.M., we have
G12  =  b G2      …...(1)     &    G22  =  c G      …...........(2)
Multiply (1) by (2)
=>G1G22 = b Gc G1   => GG= bc  or  G= bc / G2       …......(3)
Divide (1) by (2)
=> G12 / G2 =  b G2 / c G =>  c G13  = b G2 
G1= b G2/ c      …....(4) 
From (3),  G1= (bc / G2) =  bc3 / G2     ........(5)
Equate (4) & (5).    b G2/ c = bc/ G2
(G23) = b2 c4  =>  G2= b c2 
From (5), G1 = bc3 / G2= bc3 / b c
\therefore G13  = b2 c
G1+ G23  = b2 c + b c2  = bc (b + c)            ….........(6)
Also, A is the A.M. of b and c. => A = (b + c)/2  => 2A = b + c 
Substitute value of (b + c) in (6).
\therefore G1+ G2 =  bc (2A)
Thus, 
G1+ G2 = 2Abc
If satisfied with the solution, pls approve.